The step-down chopper has the following datar Ra=0:35 E-480V, and Eb=20 V. Old=190A and chopper frequency=300 Hz Assuming the load current waveform to be linear, calculate the inductance L. which would limit the maximum load current ripple to 14% of the average load current


1
Expert's answer
2021-02-10T01:11:31-0500

"\Delta L = \frac{Vdc(1-D)^2Dt}{L}"

"8 = \frac{480(1-0.04167)^20.4167}{L\times 300}"

"L= 7.654 mH"